A charge q is placed at point D of the cube. Find the elctric flux passing through the face EFGH and AEHD.


Grade:12

1 Answers

pratham ashish
14 years ago
Hi as w know from gauss law flux thruogh the cube if charge is placed on one of its corner = q/8(epsilon) but if u consider the feces of cube ABCD ,ADEH ,HDCG these all r perpendicular to the charge placed so electric field lines will b perpendicular to these faces so all the flux will b passng thru the faces ABFE , EFGH , & BCFG AND allthese faces are symmmtrical to charge so , equal amount of flux will b pasing thru all these i.e {q/8(epsilon)}/3 = q/24(epsilon) (form every of three faces mentioned above)